OpenStudy (bballanic):

Solve for x. x/6 - x/9 = 9/7

OpenStudy (astrophysics):

Ignore the variable for a second, what is \[\frac{ 1 }{ 6 }-\frac{ 1 }{ 9 }?\]

OpenStudy (bballanic):

1/18 ?

OpenStudy (photon336):

Yeah, actually you can re-write it \[\frac{ x }{ 6 }-\frac{ x }{ 9 } = x(\frac{ 1}{ 6 }-\frac{ 1 }{ 9 })\] so first \[\frac{ 1 }{ 6}-\frac{ 1 }{ 7 }\]

OpenStudy (photon336):

sorry the 7 should not be there

OpenStudy (astrophysics):

Yes exactly, so we have \[\frac{ x }{ 18 } = \frac{ 9 }{ 7 }\]

OpenStudy (astrophysics):

Can you solve for x now?

OpenStudy (bballanic):

23.14?

OpenStudy (astrophysics):

Yes that's right but we like things to be exact in math, \[x = \frac{ 18(9) }{ 7 } = \frac{ 162 }{ 7 }\]

OpenStudy (bballanic):

oh ok! thanks so much! :)

OpenStudy (astrophysics):

Anytime!

OpenStudy (bballanic):

apparently the answer was 54

OpenStudy (astrophysics):

Are you sure you put the question correctly? \[\frac{ x }{ 6 }-\frac{ x }{ 9 }=\frac{ 9 }{ 7 }\]

OpenStudy (astrophysics):

Because that is 162/7